OpenStudy (anonymous):

What are the solutions of the system? Use a graphing calculator. (i don't have a graph calc.) y=x^2+4.6+4.7 y=3.3x-3.7

OpenStudy (radar):

X^2+9.3=3.3x-3.7 x^2-3.3x+13=0 solve as a quadratic. The first equation has no x terms, did you copy it correctly?

OpenStudy (anonymous):

the first one was y=x^2+4.6x+4.7

OpenStudy (radar):

Thought that might be case, that changes things: y=y x^2+4.6x+4.7=3.3x-3.7 x^2 + 1.3x+ 8.4=0 Solve that as quadratic using quadratic formula

OpenStudy (radar):

The roots would be using the quadratic formula\[-1.3\pm \sqrt{1.3^{2}-4(8.4)}\over2\]

OpenStudy (radar):

\[-1.3\pm \sqrt{1.69-33.6}\over2\]\[-1.3\pm \sqrt{-31.91}\over2\]

OpenStudy (radar):

Both roots are complex, nor real roots.

OpenStudy (anonymous):

okay thank you!!!